5. An automobile goes around a slippery curve of radius 52 m. If the road is banked at an angle of 18 with the horizontal, at what maximum speed in kph may the automobile travel to avoid the tendency to skid? 45.892 kph 44.120 kph 43.120 kph 46.346 kph
Answer #1

Maximum speed on a banked road to avoid skid is given by:

v = sqrt(gR tan(th)

= sqrt (9.82*52*tan18)

= 12.87 m/s

= 12.87431689 * 3.6 kph

= 46.346 kph